OpenStudy (anonymous):

what is the value of x in the equaion 2(x-4)=4(2x+1) ?

OpenStudy (anonymous):

\[2(x-4)=4(2x+1)\] \[2x-8=8x+4\] \[-8-4=8x-2x\] \[-12=6x\] \[x=\frac{-12}{2}=-6\]

OpenStudy (anonymous):

the only options are -2 ,2, -1\2, 1\2 ...